An electric heater has a resistance of R = 50ohms. It is connected across a power
supply of V = 240 V. for a period of T, I = 2 hr. Determine the work done (W) by
the power supply on the heater, How many ‘units' of electricity are consumed?
1
Expert's answer
2020-10-01T09:57:53-0400
  • "Energy=power\times time"


"=\frac{V^2}{R}\times 2 hours"


"=\frac{240^2}{50}\times 2=2304Wh"


Work done/ Energy= 2304Wh


  • Units of electricity consumed


"=\frac{2304}{1000}=2.304KWH"

=2.304 units